# Kummer-Artin-Schreier-Witt Theory

### v1

@@ 1 Introduction
Our motivation is the lifting problem for Galois covers of curves, which asks: given a finite group $G$ and a $G$-cover $\phi$ of a curve in characteristic $p$, does $\phi$ lift to characteristic 00? It is important to note that, in general, the answer is NO. As with many problems in Galois theory, it is natural to begin by investigating the case where $G$ is abelian. There have been many breakthroughs recently; for instance, Obus, Wewers, and Pop have proved that all cyclic covers lift [8, 9].
Recall that every abelian cover can be realized as the pullback of a suitable isogeny of algebraic groups [12]. Therefore, to study the lifting problem for abelian covers, one can instead explore the lift of cyclic isogenies of group schemes from characteristic $p$ to characteristic 00. A classical argument allows us to further restrict to the case where $G$ is cyclic of order $p^{s}$ and the isogeny Weierstrass-$\wp$ of the ring $W_{s}$ of Witt vectors of length $s$, which gives the following exact sequence
$0\xrightarrow{}\mathbb{Z}/p^{s}\xrightarrow{}W_{s}\xrightarrow{\wp}W_{s}\rightarrow 0$ (1)
of group schemes over $\overline{k}$, where ${\rm char}(k)=p$. Sekiguchi and Suwa showed that (1) lifts .
@@
of group schemes. The generic fiber of (2) is isomorphic to the Kummer exact sequence
@@ 1.1 Outline
In Section 2, we recall Kummer theory and Artin-Schreier-Witt theory, which classify tamely ramified and wildly ramified cyclic extensions, respectively. In the same section, we restate Matsuda’s theory, which distinguishes our method from the one used by Sekiguchi and Suwa. Section 3 defines the cover on the level of field extension. More precisely, the section constructs the Kummer class of the extension, which is basically the $G_{i}$’s from (5) by truncating certain power series that arise from Matsuda’s theory. Section 4 derives the $E_{i}$’s from (4) using the Kummer extension previously defined. Section 5 shows that our previous constructions give a cyclic étale cover of an affine scheme with goood reduction. Section 6 further proves that the cover is an isogeny. Finally, Section 7 gives the proof of Theorem 1.2.
@@ 1.2 Notations
The letter $K$ will always denote a field of mixed characteristic $(0,p)$ with uniformizer $\pi$ and a discrete valuation $\nu:K^{\times}\rightarrow\mathbb{Q}$, normalized so that $\nu(p)=1$. The ring of integers of $K$ will be denoted by $R$, and its residue field $k=R/(\pi)$ is perfect. An example to keep in mind is $K=\mathbb{Q}_{p}[\zeta_{p^{s}}]$, with $\pi=\zeta_{p^{s}}-1$, which has valuation $\nu(\pi)=\frac{1}{p^{s-1}(p-1)}$. Set $\lambda=\zeta_{p}-1$, a constant that will appear frequently.
We fix an algebraic closure $\overline{K}$ of $K$. The symbol $\mathbb{K}$ usually denotes a Laurent series over $K$. The valuation $\nu$ extends to the Gauss valuation on $\mathbb{K}$. For instance, if $\mathbb{K}=K((X))$ and $f(X)=\sum_{i\geq n}a_{i}X^{i}\in\mathbb{K}$, then $\nu(f)=\inf_{i\geq n}\{\nu(a_{i})\}$.
@@ 1.3 Acknowledgements
We thank Mézard and Tossici for useful conversations. The first author thanks the National Center for Theoretical Sciences, National Taiwan University for the excellent working conditions.
@@ 2.1 Artin-Schreier-Witt theory
$\begin{split}0\xrightarrow{}\mathbb{Z}/p^{s}\cong W_{s}(\mathbb{F}_{p})\xrightarrow{}W_{s}&\xrightarrow{\wp}W_{s}\rightarrow 0,\\ \underline{x}&\mapsto F\underline{x}-\underline{x}\end{split}$ (7)
of group schemes over $\overline{k}$, which induces the identification
@@ 2.2 Kummer Theory
Thus, each $\chi\in\operatorname{H}^{1}(K,\mathbb{Z}/p^{s}\mathbb{Z})$ can be associated with a rational function $G$ in $K^{\times}/(K^{\times})^{p^{s}}$. Additionally, the function $G$ can be factored as follows:
@@
$(K^{\times})^{\oplus s}/\phi_{s}((K^{\times})^{\oplus s})$ ${\sim}K^{\times}/(K^{\times})^{p^{s}}\xrightarrow[]{\sim}\operatorname{H}^{1}(K,\mathbb{Z}/p^{s}\mathbb{Z})$
$(a_{1},\ldots,a_{s})$ $\mapsto[a_{1}a_{2}^{p}\ldots a_{s}^{p^{s-1}}]\mapsto\mathfrak{K}_{s}(a_{1}a_{2}^{p}\ldots a_{s}^{p^{s-1}}),$
@@
$(K^{\times})^{\oplus s}/\phi_{s}((K^{\times})^{\oplus s})$ is a group under coordinate-wise multiplication. We say $(a_{1},\ldots,a_{s})\sim(b_{1},\ldots,b_{s})$ if there exists $(c_{1},\ldots,c_{s})$ such that
@@ 2.3.2 Matsuda’s Kummer-Artin-Schreier-Witt theory
The ring $\mathcal{R}^{\operatorname{int}}$ is a Henselian discrete valuation ring. Consequently, $\mathcal{R}^{\rm bd}$ is a Henselian local field with residue field $\kappa:=k((z))$.
@@
Remark 2.8.
We believe Matsuda’s theory has potential applications far beyond the scope of this paper. For instance, expanding the domain and the range of (14) enables us to address inseparable extensions. This will be explored in our future work on the refined Swan conductor [2], an important invariant that measures how far a wildly ramified cover is from having separable reduction.
@@ 3.2 Construct the Kummer Extension
Let
$f=\sum_{n_{0},\ldots,n_{s}\geq 0}a_{n_{0},\ldots,n_{s}}X_{0}^{n_{0}}\cdots X_{s}^{n_{s}}\in R\llbracket X_{0},\ldots,X_{s}\rrbracket$
satisfy the Tate condition if, for every multi-index $n=(n_{0},\ldots,n_{s})$, the coefficients $a_{n}$ satisfy:
$\nu(a_{n})\to\infty\quad\text{as}\quad|n|=n_{0}+\cdots+n_{s}\to\infty.$
@@
Next, we define $G_{s}\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}][X_{0},\ldots,X_{s-2}]$ to be the polynomial obtained by truncating $G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))$ at terms with valuations at most $\frac{p}{p-1}$. Assertion 2 is immediately verified. Assertion 3 can be derived from the computation of $X_{s-1}^{\prime}$ in the proof of Proposition 3.3. Recall that
@@ 3.3 Defining the étale Cover
$Z_{s-1}^{p^{s}}=\prod_{i=0}^{s-1}(G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i})^{p^{i}},$ (17)
where $G_{1}=1$.
Due to Theorem 2.7 and Proposition 3.3, the special fiber of $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is birational to the Artin-Schreier-Witt extension of $\operatorname{Proj}k[x_{0},\ldots,x_{s-1},y]$, where $x_{i}$ (resp. $y$) is the reduction of $X_{i}$ (resp. $Y$) modulo $\pi$, given by the Witt vector $(x_{0},\ldots,x_{s-1})$.
@@ 4.2 The Case $s>1$
$Y_{i-1}^{\prime}=(-1)^{p}\frac{\lambda^{p-1}}{p}Y_{i-1}^{p}+Y_{i-1}+F_{i}(Y_{0},\ldots,Y_{i-1}),$
@@
4.3 Approximation of the Power Series $E_{s,p}$
@@
where $\tilde{X}_{j}=X_{j-1}^{p}+pR$. By the induction hypothesis, $\nu(E_{i-1,p}((\tilde{X}_{1},\ldots,\tilde{X}_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}$, which is strictly less than 1 in general, except when $p=2$ and $i=2$. Thus, $\nu(E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}$ for $p>2$ or $p=2$ and $i>2$. A direct computation confirms that $\nu(E_{2,2}((X_{0},X_{1}))-1)=\frac{1}{2}$.
@@
Since $\nu(E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}<1$ unless $p=2$ and $i=2$, we deduce that $\nu(A^{p}+\sum_{j=1}^{p-1}\binom{p}{j}A^{j})=p\nu(A)$. Therefore, we conclude that $\nu(E_{i,p}((X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-1}(p-1)}$, completing the proof. ∎
@@
$\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+g_{s-1}(Y_{0},\ldots,Y_{s-1})-(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})X_{s-1}=0.$ (25)
@@ 5.1 Case $s=1$
$R\left[Y_{0},\frac{1}{1-\lambda Y_{0}}\right]$ is the integral closure of $R\left[X_{0},\frac{1}{1+p\lambda X_{0}}\right]$ in $L_{1}$.
@@
Thus, $\mathcal{W}_{1}=\operatorname{Spec}R\left[Y_{0},\frac{1}{1-\lambda Y_{0}}\right]$ has good reduction $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]$. Additionally, from (27) and Kummer theory, the cover $\mathcal{Y}_{1}\rightarrow\mathbb{P}^{1}_{R}$ branches generically at $X_{0}=\infty$ and $X_{0}=\frac{1}{p\lambda}$. Therefore, $\mathcal{W}_{1}\rightarrow\mathcal{V}_{1}$ is étale with Galois group $\mathbb{Z}/p$, and its special fiber is isomorphic to the Artin-Schreier extension $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]\rightarrow\operatorname{Spec}\mathbb{F}_{p}[x_{0}]$ given by $-y_{0}^{p}+y_{0}=x_{0}$.

### v2

@@ 1 Introduction
Our motivation is the lifting problem for Galois covers of curves, which asks: given a finite group $G$ and a $G$-cover $\phi$ of a curve in characteristic $p$, does $\phi$ lift to characteristic 00? It is important to note that, in general, the answer is NO. As with many problems in Galois theory, it is natural to begin by investigating the case where $G$ is abelian. There have been significant breakthroughs recently; for instance, a joint effort by Obus, Wewers, and Pop has demonstrated that all cyclic covers of (smooth projective) curves lift [8, 9].
Recall that every abelian cover can be realized as the pullback of a suitable isogeny of algebraic groups [12]. Therefore, to study the lifting problem for abelian covers, one can instead explore the lift of cyclic isogenies from characteristic $p$ to characteristic 00. A classical argument allows us to further restrict to the case where $G$ is cyclic of order $p^{s}$ and the isogeny Weierstrass-$\wp$ of the ring $W_{s}$ of Witt vectors of length $s$, which gives the following exact sequence:
$0\xrightarrow{}\mathbb{Z}/p^{s}\xrightarrow{}W_{s}\xrightarrow{\wp}W_{s}\rightarrow 0,$ (1)
of group schemes over $\overline{k}$, where ${\rm char}(k)=p$. Sekiguchi and Suwa claimed that (1) lifts.
@@
of group schemes over $\overline{R}$. The generic fiber of (2) is isomorphic to the Kummer exact sequence
@@
Remark 1.4.
We believe that Matsuda’s theory has potential applications far beyond the scope of this paper. For instance, expanding the domain and range of (6) may enable us to address inseparable extensions. This will be explored in our future work on the refined Swan conductor introduced in [2], an important invariant that measure how far a wildly ramified cover is from having separable reduction.
@@ 1.1 Outline
In Section 2, we recall Kummer theory and Artin-Schreier-Witt theory, which classify tamely ramified and wildly ramified cyclic extensions, respectively. In the same section, we restate Matsuda’s theory, which distinguishes our method from that used by Sekiguchi and Suwa. Section 3 defines the cover at the level of field extensions. More precisely, the section constructs the Kummer class of the extension, which corresponds to the $G_{i}$’s in (5), obtained by truncating certain power series arising from Matsuda’s theory. Section 4 derives the $E_{i}$’s in (4) using the previously defined Kummer extension. Section 5 demonstrates that our constructions yield a cyclic étale cover of an affine scheme with good reduction. Section 6 further proves that the cover is an isogeny, establishing Theorem 1.1. Finally, Section 7 provides the proof of Theorem 1.2.
@@ 1.2 Notations
The letter $K$ will always denote a field of mixed characteristic $(0,p)$ with uniformizer $\pi$ and a discrete valuation $\nu:K^{\times}\rightarrow\mathbb{Q}$, normalized so that $\nu(p)=1$. The ring of integers of $K$ will be denoted by $R$, and its residue field $k=R/(\pi)$ is perfect. An example to keep in mind is $K=\mathbb{Q}[\zeta_{p^{s}}]$, with $\pi=\zeta_{p^{s}}-1$, which has valuation $\nu(\pi)=\frac{1}{p^{s-1}(p-1)}$. Set $\lambda=\zeta_{p}-1$, a constant that will appear frequently.
We fix an algebraic closure $\overline{K}$ of $K$. The symbol $\mathbb{K}$ usually denotes a Laurent series over $K$. The valuation $\nu$ extends to the Gauss valuation on $\mathbb{K}$. For instance, if $\mathbb{K}=K\mathopen{(\!(}X\mathclose{)\!)}$ and $f(X)=\sum_{i\geq n}a_{i}X^{i}\in\mathbb{K}$, then $\nu(f)=\inf_{i\geq n}\{\nu(a_{i})\}$.
@@ 1.3 Acknowledgements
We thank Kiran Kedlaya, Ariane Mézard, Andrew Obus, Frans Oort, Matthieu Romagny, and Dajano Tossici for their valuable discussions. The first author expresses gratitude to the National Center for Theoretical Sciences at National Taiwan University for providing excellent working conditions.
@@ 2.1 Artin-Schreier-Witt theory
$\begin{split}0\xrightarrow{}\mathbb{Z}/p^{s}\cong W_{s}(\mathbb{F}_{p})\xrightarrow{}W_{s}&\xrightarrow{\wp}W_{s}\rightarrow 0,\\ \underline{y}=(y_{0},\ldots,y_{s-1})&\mapsto\underline{y}-\sigma(\underline{y}),\end{split}$ (7)
of group schemes over $\overline{k}$, where $\sigma$ is the Frobenius operator on Witt vectors. This induces the identification
@@
Additionally, if we write
$\wp(y_{0},\ldots,y_{s-1})=(y_{0}-y_{0}^{p},y_{1}-y_{1}^{p}+f_{2}(y_{0}),\ldots,y_{s-1}-y_{s-1}^{p}+f_{s}(y_{0},\ldots,y_{s-2})),$
where $f_{i}(y_{0},\ldots,y_{i-2})\in\mathbb{F}_{p}(y_{0},\ldots,y_{i-2})$ for $i=2,\ldots,s$ and $f_{1}=0$, then $\overline{\chi}$ is defined by a system of Artin-Schreier extensions
$y_{i-1}-y_{i-1}^{p}+f_{i}(y_{0},\ldots,y_{i-2})=x_{i-1}.$ (8)
@@ 2.2 Kummer Theory
Thus, each $\chi\in\operatorname{H}^{1}(K,\mathbb{Z}/p^{s}\mathbb{Z})$ can be associated with a rational function $G$ in $K^{\times}/(K^{\times})^{p^{s}}$. Additionally, we may assume that the function $G$ is of the following form:
@@
$(K^{\times})^{\oplus s}/\phi_{s}\big{(}(K^{\times})^{\oplus s}\big{)}$ $\xrightarrow{\sim}$ $\;K^{\times}/\big{(}K^{\times}\big{)}^{p^{s}}$ $\xrightarrow{\sim}\operatorname{H}^{1}\big{(}K,\mathbb{Z}/p^{s}\mathbb{Z}\big{)}$ (9)
$(a_{1},\ldots,a_{s})$ $\mapsto$ $\;[a_{1}a_{2}^{p}\cdots a_{s}^{p^{s-1}}]$ $\mapsto\mathfrak{K}_{s}\big{(}a_{1}a_{2}^{p}\cdots a_{s}^{p^{s-1}}\big{)},$
@@
$(K^{\times})^{\oplus s}/\phi_{s}((K^{\times})^{\oplus s})$ is a group under coordinate-wise multiplication. We say $(a_{1},\ldots\allowbreak\ldots,a_{s})\sim(b_{1},\ldots,b_{s})$ if there exists $(c_{1},\ldots,c_{s})$ such that
@@ 2.3.2 Matsuda’s Kummer-Artin-Schreier-Witt theory
The ring $\mathcal{R}^{\operatorname{int}}$ is a Henselian discrete valuation ring. Consequently, $\mathcal{R}^{\rm bd}$ is a Henselian local field with residue field $\kappa:=k\mathopen{(\!(}z\mathclose{)\!)}$.
@@ 3.2 Construct the Kummer Extension
We say that
$f=\sum_{n_{0},\ldots,n_{s-1}\geq 0}a_{n_{0},\ldots,n_{s-1}}X_{0}^{n_{0}}\cdots X_{s-1}^{n_{s-1}}\in R\llbracket X_{0},\ldots,X_{s-1}\rrbracket$
satisfies the Tate condition if, for every multi-index $n=(n_{0},\ldots,n_{s-1})$, the coefficients $a_{n}$ satisfy:
$\nu(a_{n})\to\infty\quad\text{as}\quad|n|=n_{0}+\cdots+n_{s-1}\to\infty.$
@@
Next, we obtain $G_{s}\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}][X_{0},\ldots,X_{s-2}]$ by truncating $G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))$ at terms with valuations at most $\frac{p}{p-1}$. Assertion 2 is immediately verified. Assertion 3 can be derived from the computation of $X_{s-1}^{\prime}$ in the proof of Proposition 3.3. Recall that
@@ 3.3 Defining the étale Cover
$Z_{s-1}^{p^{s}}=(1+p\lambda X_{0})\prod_{i=1}^{s-1}(G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i})^{p^{i}}.$ (19)
Proposition 3.7.
The special fiber of $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is birational to the Artin-Schreier-Witt extension of $\operatorname{Proj}k[x_{0},\ldots,x_{s-1},y]$, where $x_{i}$ (resp. $y$) is the reduction of $X_{i}$ (resp. $Y$) modulo $\pi$, given by the Witt vector $(x_{0},\ldots,x_{s-1})$.
Proof.
Applying Lemma 2.6 to $K:=\mathbb{Q}[\zeta_{p^{s}}]\mathopen{(\!(}X_{0},\ldots,X_{s-2}\mathclose{)\!)}$, which is discretely valued with respect to the Gauss valuation, and setting $z=X_{s-1}$, we obtain a Henselian ring $\mathcal{R}^{\operatorname{int}}$ and a Henselian field $\mathcal{R}^{\text{bd}}$ containing $\mathbb{Q}[\zeta_{p^{s}}](X_{0},\ldots,X_{s-1})$ with residue field $\kappa=\mathbb{F}_{p}\mathopen{(\!(}x_{0},\ldots,x_{s-1}\mathclose{)\!)}$.
It follows from the proof of Proposition 3.3 that $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})$ lies in $W_{s}(\mathcal{R}^{\operatorname{int}})$. Therefore, by Theorem 2.7, the extension of $\mathcal{R}^{\text{bd}}$ given by (19) corresponds to the Witt vector $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})$, which is a lift of $(x_{0},\ldots,x_{s-1})\in W_{s}(\kappa)$. This completes the proof. ∎
@@ 4.2 Inductive Step
$Y_{i-1}^{\prime}=(-1)^{p-1}\frac{\lambda^{p-1}}{p}Y_{i-1}^{p}+Y_{i-1}+F_{i}(Y_{0},\ldots,Y_{i-1}),$
@@
4.2.1 Approximation of the Power Series $E_{s,p}$
@@
where $\tilde{X}_{j}=X_{j-1}^{p}+pR$. By the induction hypothesis, $\nu(E_{i-1,p}((\tilde{X}_{1},\ldots,\tilde{X}_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}$, which is strictly less than 1 in general, except when $p=2$ and $i=2$. Thus, $\nu(E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}$ for $p>2$ or $p=2$ and $i>2$.
@@
Since $\nu(E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}<1$ unless $p=2$ and $i=2$, we deduce that $\nu(A^{p}+\sum_{j=1}^{p-1}\binom{p}{j}A^{j})=p\nu(A)$. Therefore, we conclude that $\nu(E_{i,p}((X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-1}(p-1)}$.
Finally, a direct computation confirms that $\nu(E_{2,2}((X_{0},X_{1}))-1)=\frac{1}{2}$, completing the proof. ∎
@@ 4.2.2 Computing the Minimal Polynomial
$\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+H_{s}(Y_{0},\ldots,Y_{s-1})-(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})X_{s-1}=0.$ (27)
@@
In fact, we can provide further details on the reduction of $F_{s}$ and $H_{s}$ in (27).
Corollary 4.13.
We have, for $j=1,\ldots,s$,
$H_{j}\equiv F_{j}\equiv f_{j}(y_{0},y_{1}+b_{1},\ldots,y_{j-2}+b_{j-2})+\wp(b_{j-1})\pmod{(\zeta_{p^{j}}-1)},$
where $f_{j}$ is the polynomial in (8) and $b_{j-1}\in\mathbb{F}_{p}[y_{0},\ldots,y_{j-2}]$. In particular, the special fiber of $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is birational to the cover of $\operatorname{Proj}k[x_{0},\ldots,x_{s-1},y]$ defined by
$\wp(y_{0},y_{1}+b_{1},\ldots,y_{s-1}+b_{s-1})=(x_{0},x_{1},\ldots,x_{s-1}).$
Proof.
Due to §4.1, the corollary holds for $s=1$.
For $s>1$, recall that the special fiber of $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is birational to the extension of $\mathbb{F}_{p}[x_{0},\ldots,x_{s-1}]$ associated with the Witt vector $(x_{0},\ldots,x_{s-1})$, which can be defined as a system of Artin-Schreier extensions as in (8). By induction, the last of these extensions is given by
$y_{s-1}-y_{s-1}^{p}+f_{s}(y_{0},y_{1}+b_{1},\ldots,y_{s-2}+b_{s-2})+\wp(b_{s-1})=x_{s-1},$
for some $b_{s-1}\in\mathbb{F}_{p}[y_{0},\ldots,y_{s-2},x_{s-1}]$. On the other hand, by (27), we have
$y_{s-1}-y_{s-1}^{p}+\overline{H}_{s}=x_{s-1},$
where $\overline{H}_{s}$ denotes the reduction of $H_{s}$ modulo $(\zeta_{p^{s}}-1)$. The result follows immediately. ∎
Example 4.14.
When $p=2$, $\overline{H}_{2}=y_{0}^{4}+y_{0}^{8}=0\wp(y_{0})^{4}$ if we choose $G_{2}(Y_{0})=1+2\zeta_{4}Y_{0}-52Y_{0}^{4}$. This is the lowest degree polynomial for $G_{2}$ that we can select.
@@ 5.1 The Case $s=1$
$R\left[Y_{0},\frac{1}{1+\lambda Y_{0}}\right]$ is the integral closure of $R\left[X_{0},\frac{1}{1+p\lambda X_{0}}\right]$ in $L_{1}$.
@@
Thus, $\mathcal{W}_{1}=\operatorname{Spec}R\left[Y_{0},\frac{1}{1+\lambda Y_{0}}\right]$ has good reduction $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]$. Additionally, from (29) and Kummer theory, the cover $\mathcal{Y}_{1}\rightarrow\mathbb{P}^{1}_{R}$ branches generically at $X_{0}=\infty$ and $X_{0}=\frac{1}{p\lambda}$. Therefore, $\mathcal{W}_{1}\rightarrow\mathcal{V}_{1}$ is étale with Galois group $\mathbb{Z}/p$, and its special fiber is isomorphic to the Artin-Schreier extension $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]\rightarrow\operatorname{Spec}\mathbb{F}_{p}[x_{0}]$ given by $-y_{0}^{p}+y_{0}=x_{0}$.